OpenStudy (anonymous):

A bricklayer uses a cylinder to store sand. The cylinder is 5 ft tall and has a radius of 0.5 ft. Each batch of mortar contains 0.48 ft3 of sand. How many batches of mortar can the bricklayer make with 1 cylinder of sand? Use 3.14 to approximate pi, andround your final answer to the nearest whole number. @paki @iambatman

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

i will medal

OpenStudy (welshfella):

first find the volume of the cylinder this = 3.14*r^2*h where r = radius and h = height

OpenStudy (anonymous):

ok so what would be V=3.14r^2H

OpenStudy (welshfella):

then to find the number of batches of mortar divide this volume by 0.48

OpenStudy (anonymous):

ok

OpenStudy (welshfella):

V = 3.14 * 0.5^2 * 5 ft3

OpenStudy (anonymous):

3.925 FT3

OpenStudy (welshfella):

right

OpenStudy (anonymous):

so wut now

OpenStudy (welshfella):

so no. of batches = 3.925 / 0.48

OpenStudy (anonymous):

8.177083333

OpenStudy (anonymous):

??? ><__><

OpenStudy (welshfella):

right again - round this to nearest whole numb er = ?

OpenStudy (anonymous):

8

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

so dats my answer?

OpenStudy (anonymous):

@welshfella Is that the answer?

OpenStudy (anonymous):

I GOT IT RIGHT...Thx @welshfella

OpenStudy (welshfella):

yw